OpenStudy (anonymous):

y = -2x – 6 Find the x- and y-intercepts for the equation. Please show all of your work

OpenStudy (mertsj):

Find the x intercept by letting y = 0. Find the y intercept by letting x = 0

OpenStudy (callisto):

y = -2x – 6 2x +y +6 =0 for a equation having general form : ax +by+c =0 , where a, b, c are constants slope = -a/b x-int = -c/a y-int = -c/b if it is in a form : y = mx + c where m, c are constants m= slope c= y- int of course the easiest way to find x-int is to put y=0 and to find y-int, just put x=0 try to solve it and you'll get the answers

OpenStudy (anonymous):

how would you plug in the 0

OpenStudy (mertsj):